OpenStudy (anonymous):

Multiply and simplify. Assume that all variable represent nonnegative numbers. Help with this problem please.

OpenStudy (anonymous):

\[(-4\sqrt{5})(-2\sqrt{3})\]

OpenStudy (anonymous):

we have : \[\sqrt{15}\]

OpenStudy (anonymous):

and (-2)*(-4)

OpenStudy (anonymous):

how do I sovle it

OpenStudy (anonymous):

We have : \[\sqrt{15}*8\]

OpenStudy (anonymous):

@tashasp :Got it ?:)

OpenStudy (anonymous):

thinks so

OpenStudy (anonymous):

I multiply the square root 15*8

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\[(-4\sqrt{5})(-2\sqrt{3})\]

OpenStudy (anonymous):

help please